A college graduate in 1972 found a job paying $7,200. The CPI was 0.418 in 1972. A college graduate in 2018 found a job paying $50,390. The CPI was 0.2511 in 2018.

Calculate real income for 1972
Calculate real income for 2018
Calculate growth rate of nominal income between 1972 and 2018
Calculate growth rate of real income between 1972 and 2018
1
Expert's answer
2020-04-17T09:43:18-0400

"RealIncome=\frac{Wages}{1+Inflation Rate}"

Inflation is equal to CPI.

"RI(1972)=\frac{7,200}{1+0.418}=\$5,077.74"

"RI(2018)=\frac{50,390}{1+0.2511}=\$40,276.56"

"GrowthRate=\frac{present}{past}^{1/n}-1"

"GrowthRate(NI)=\frac{50,390}{7,200}^{1/46}-1=4.3%" %

"GrowthRate(NI)=\frac{40,276.56}{5,077.574}^{1/46}-1=4.6%" %